To treat a burn on his hand, a person decides to place an ice cube on the burned skin. The mass of the ice cube is 13.8 g, and its initial temperature is −12.3 ∘C. The water resulting from the melted ice reaches the temperature of his skin, 30.8 ∘C. How much heat is absorbed by the ice cube and resulting water? Assume that all of the water remains in the hand.
1
Expert's answer
2019-07-18T07:52:09-0400

Q1= mC∆T = 13.8g*2.087 J/g/deg*12.3 ∘C = 354.2 J

Q2 = m∆Hf = 13.8g*333.6 J/g = 4603.7 J

Q3= mC∆T = 13.8g*4.184 J/g/deg*30.8 ∘C = 1778.4 J

Q=1778.4+4603.7+354.2=6736.3J

Answer:6736.3J